# Certifiable Deep Learning for Reachability Using a New Lipschitz Continuous Value Function

### v2

@@ VII-B Hypothesis 2: Our online RA set certification methods can be computed in real-time
Figure 8: The convergence of the critic loss of max-min DDPG under our new Bellman equation (Ours) and under the approximated Bellman equation (Baseline) from prior works [15, 12], with different $\gamma$ values. Our critic loss under $\gamma=0.95$ converges more rapidly than under $\gamma=0.9999$ possibly due to the Lipschitz continuity of ${V_{\gamma}(x)}$ when $\gamma=0.95$. Both the baseline and ours are trained using the same training parameters and random initializations. For drone racing, we parameterize each of ${\pi_{\theta_{u}}}$, ${\phi_{\theta_{d}}}$ and ${Q_{\theta_{q}}}$ using a 4-layer ReLU NN, with 512 neurons in each layer. For the highway example, we parameterize each ${\pi_{\theta_{u}}}$, ${\phi_{\theta_{d}}}$ and ${Q_{\theta_{q}}}$ using a 3-layer ReLU NN, with 512 neurons in each layer. The NN training is conducted under 5 different random seeds.
Figure 9: The Lipschitz continuity of our new RA value functions appears to accelerate learning, and this phenomenon is not limited to max-min DDPG. In the first subfigure, we apply max-min DDPG under our new Bellman equation (Ours) and the approximated Bellman equation (Baseline) from prior works [12, 15]. In the second subfigure, we apply max-min SAC [12, 15] under both Bellman equations. The dynamics considered are the same as in Figure 2, with $r(x)=\max(\min(2-x^{2},10),-10)$ and $c(x)=\max(\min(x+1.5,10),-10)$. Due to the simplicity of this problem, we model ${Q_{\theta_{q}}}$ as a 2-layer ReLU NN with 16 neurons in each layer, and model both ${\pi_{\theta_{u}}}$ and ${\phi_{\theta_{d}}}$ as 2-layer ReLU NNs, each with 4 neurons per layer. We train the NNs under 5 different random seeds. Under $\gamma=0.95$, ${V_{\gamma}(x)}$ is Lipschitz continuous and our method converges rapidly. However, as $\gamma\to 1$, our learning process becomes unstable possibly due to the loss of Lipschitz continuity of ${V_{\gamma}(x)}$.
@@
VII-C Hypothesis 3: The Lipschitz continuity of our new value function appears to accelerate learning
@@
In Figures 8 and 9, we compare the critic loss (defined in (15) to measure the Bellman equation error) under different Bellman equations and time-discount factor $\gamma$ values. We observe that the critic loss under our new Bellman equation converges rapidly when $\gamma$ is chosen to ensure the Lispchitz continuity of our new value function.
VII-D The trade-off of selecting a time-discount factor $\gamma$
@@
We summarize our result in Figures 6 and 7. With a small $\gamma$, the learned RA sets can be conservative due to numerical errors, as an initial state whose optimal trajectory reaches the target set at a later stage may have near-zero or even negative time-discounted RA measures. However, the optimal policy tends to drive the state to the target set rapidly, as depicted in Figure 7. Conversely, when $\gamma$ is high, the value function could be discontinuous, leading to unstable learning and resulting in a suboptimal value function. In the drone racing and highway experiments, we find that $\gamma=0.95$ ensures the Lipschitz continuity of ${V_{\gamma}(x)}$, thereby enhancing learning efficiency, and also mitigates unnecessary conservatism.
VIII Conclusion
@@
We propose a new framework for learning trustworthy RA sets. Our method features a newly designed time-discounted RA value function, which is provably Lipschitz continuous and has a Bellman operator that is a contraction mapping. This enhances computational efficiency by eliminating the need for the time-discount factor annealing process used in previous works. We use max-min DDPG to learn our value functions and propose two efficient methods to certify if a set of states can safely reach the target set despite any disturbance, with deterministic guarantees. We validate our methods in drone racing hardware experiments and highway take-over simulations. In future research, we plan to extend this work to online reachability learning in uncertain environments and propose more efficient and less conservative reachability set certification methods by leveraging the problem structure.
IX Acknowledgement
We appreciate Andrea Bajcsy and David Fridovich-Keil for valuable discussion.

### v5

@@ VIII-B Hypothesis 2: Our online RA set certification methods can be computed in real-time
VIII-C Hypothesis 3: The Lipschitz continuity of our new value function appears to accelerate learning
@@
In Figure 6, we compare the critic loss (defined in (15) to measure the Bellman equation error) under different Bellman equations and time-discount factor $\gamma$ values. Our critic loss converges rapidly when $\gamma$ is chosen to ensure the Lipschitz continuity of our new value function.
VIII-D The trade-off of selecting a time-discount factor $\gamma$
@@
We summarize our result in Figures 7 and 8. With a small $\gamma$, the learned RA sets can be conservative due to numerical errors, as an initial state whose optimal trajectory reaches the target set at a later stage may have near-zero time-discounted RA measures. However, the optimal policy tends to drive the state to the target set rapidly, as depicted in Figure 8. Conversely, a large $\gamma$ can induce discontinuities in the value function, destabilizing learning and yielding suboptimal solutions. In the drone racing and highway experiments, we find that $\gamma=0.95$ ensures the Lipschitz continuity of ${V_{\gamma}(x)}$, thereby enhancing learning efficiency, and also mitigates unnecessary conservatism.
IX Conclusion and Future Work
@@
We propose a new framework for learning trustworthy reach-avoid (RA) sets. Our method features a newly designed RA value function that offers improved computational efficiency. We employ max-min DDPG to learn our value functions and propose two efficient methods to certify whether a set of states can safely reach the target set with deterministic guarantees. We validate our methods through drone racing hardware experiments and highway take-over simulations. Our certification methods can be performed in real time, but they rely on offline value function learning beforehand. Future research may explore online reachability learning, as well as more efficient RA set certification methods.